During field trip joined by 800 pupils, 610 took the rides provided by the school . Find p and q , where p was the population of children who took the rides

1
Expert's answer
2022-02-04T11:08:51-0500

"p=\frac{610}{800}=0.7625"


"q=1-p=1-0.7625=0.2375"